The perimeter of ABCDE is 130 cm. Find the length of segment AE.
1
Expert's answer
2020-09-23T12:19:45-0400

Perimeter = "\pi d" ;AE=Diameter


130="\frac{22}{7}" d


d="\frac{7}{22} ×130"


"d" ="\frac{910}{22}"


"d" =41.36cm